### v1

@@ 4.3 Examples, revisited
$\text{HHI}(\mathbf{y}^{\star})=\operatorname{MRCI}(D_{\mathrm{pool}};\geq_{FSD},\mathcal{F})=\Big(\tfrac{15}{16}\Big)^{2}+\Big(\tfrac{1}{16}\Big)^{2}=\tfrac{113}{128}\approx 0.883,$
so that $N_{\mathrm{eff}}\approx 1.13$. The pooled behavior is thus close to being explained by a single heuristic (salience), with MAP or an attention rule playing a purely residual role in one menu.
Stability scores and concentration gains refine this picture. Because attention rules can always rationalize any single observation, MAP is never indispensable for sustaining $H^{\star}=\operatorname{MRCI}(D_{\mathrm{pool}};\geq_{FSD},\mathcal{F})$: there are deletion orders in which MAP can be removed without changing the optimal MRCI value (the corresponding observation can be reassigned to A1 or A2). Hence $\kappa(\textsc{MAP})=0$, and in this pooled illustration $\phi(\textsc{MAP})=0$ as well.
For SAL, removing it makes maximal concentration fall. Indeed, when SAL is excluded from the library, many observations must be reassigned to REG and to the attention rules, which reduces the maximal HHI sharply. A conservative reassignment yields
$\mathrm{HHI}\approx\Big(\tfrac{10}{16}\Big)^{2}+\Big(\tfrac{6}{16}\Big)^{2}\approx 0.53,$
so the implied concentration gain of SAL is at least
$\phi(\textsc{SAL})\ \geq\ \frac{0.883-0.53}{0.883}\ \approx\ 0.4.$
Moreover, SAL appears in every deletion-stable set across permutations, so $\kappa(\textsc{SAL})=1$: it is order-robustly indispensable for sustaining the maximally concentrated explanation.
In this pooled illustration, MRCI indicates a highly parsimonious structure ($N_{\mathrm{eff}}\approx 1$), while the stability score and concentration gain reveal which rule carries this parsimony. Salience emerges as the workhorse decision rule: it is always retained in value-preserving prunings of the library and accounts for a substantial share of the attainable concentration, whereas other rules such as MAP are weakly identified from the standpoint of rule concentration.
5 Computational Strategy
An exact formulation of this problem as a mixed–integer quadratic program (MIQP) is provided in Appendix D. However, maximizing a convex objective function (concentration) over discrete assignments is an NP-hard problem. As demonstrated in our benchmarking exercise (see Appendix D), the MIQP approach scales fast with sample size, rendering it computationally intractable for standard datasets. To address this, we employ a heuristic algorithm that exploits the structure of the objective. Since we are maximizing concentration, optimal solutions by convexity tend to load on rules that can rationalize many menus simultaneously. The algorithm below is designed to efficiently explore such high–coverage, high–concentration assignments.
5.1 Overview of the heuristic
The procedure has three components:
  1. i
  2. $i$
  3. iii
We now describe each component in turn.
Phase 1: Pre–processing
Local admissibility.
$a_{t,f}:=\mathbf{1}\{f\in\mathcal{F}_{t}^{\mathrm{strict}}\},$
Phase 2: Greedy assignment for a fixed rule order
Rule ordering.
$\text{coverage}(f):=\sum_{t=1}^{T}a_{t,f},$
State variables.
Greedy pass.
  1. i
  2. $i$
Phase 3: Random restarts and selection
5.2 Accuracy and Scalability
The heuristic replaces combinatorial optimization with a sequence of fast linear passes. To validate this approach, we benchmarked the heuristic against the MIQP solver on a set of 3,430 subsampled optimization tasks (see Appendix D for full details).
@@ 6.1 Random Rule Models and two assumptions
  • A probability vector $\pi=(\pi_{f})_{f\in\mathcal{F}}\in\Delta^{|\mathcal{F}|-1}$
  • An i.i.d. latent process $(H_{t})_{t\in\mathcal{T}}$ with $\Pr(H_{t}=f)=\pi_{f}$. Given a menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$ and a realized latent rule $H_{t}=f$, the agent chooses the option prescribed by $f$ whenever $f$ yields a strict dominance recommendation at $A_{t}$; when $f$ does not strictly discriminate at $A_{t}$, choice is generated by a fixed fallback that does not depend on the realized menu index. [^9] Any fixed fallback is admissible for the interpretation below. In particular, one may take the fallback to be a randomization between A1 and A2 with fixed weights.
@@ 7 Empirical Application
We apply our framework to the publicly available CPC18 competition dataset (erev_ert_plonsky_2017; plonsky2017_cpc18). [^10] The data can be downloaded on the following website: https://cpc-18.com/data/. Since our rule library and dominance benchmark are defined over lotteries with known probabilities, we restrict attention to the risk problems and exclude ambiguity tasks. The resulting sample comprises $N=686$ subjects (average age 24.9, s.d. 2.93; 54% female) recruited across multiple locations in Israel. Each subject faces $20$ to $28$ distinct binary lottery choice problems, and each problem is repeated for $25$ trials, yielding $500$ to $700$ decisions per subject. The design includes both no-feedback and feedback regimes, [^11] In the feedback regime, repeated exposure and realized outcomes may induce learning or history-dependence. and the choice problems were design to elicit canonical behavioral anomalies in risky choice, including Allais paradoxes (allais1953; tversky_kahneman1986), certainty-like effects (kahneman_tversky1979), break-even effects (thaler_johnson1990), splitting effects (birnbaum2008), or systematic departures from expected-value reasoning driven by the weighting of rare events (barron_erev2003). For a detailed description of the experimental design, the exact list of problems, and the mapping from problems to targeted phenomena, see erev_ert_plonsky_2017. Summary statistics are provided in Table 5.
@@
  • Notes: The table reports summary statistics for the $N=686$ subjects in the CPC18 dataset. The lower panel reports the Maximum Rule Concentration Index (MRCI), computed subject-by-subject under the rule library $\mathcal{F}$ and the $\geq_{\mathrm{FSD}}$ benchmark. MRCI is estimated with the heuristic algorithm in Section 5 using $R=100$ independent random restarts per subject; the Effective Number of Rules is $N_{\mathrm{eff}}=1/\operatorname{MRCI}$. The “Share Non-Random” rows report the fraction of subjects for whom we reject the exchangeable random-choice null using the finite-sample permutation test in Section 6. We approximate the exact permutation distribution by Monte Carlo with $B=500$ random permutations. Each permuted-statistic $\operatorname{MRCI}^{(b)}$ is computed using the same heuristic (with $R=100$ restarts).
@@
For each subject $i$, we compute the Maximum Rule Concentration Index $\operatorname{MRCI}(D_{i};\,\geq_{\mathrm{FSD}},\mathcal{F})$ using the optimization routine described in Section 5. In practice, we first pre-compute the menu-by-rule admissibility matrix (which records, for each binary problem faced by the subject, which rules in $\mathcal{F}$ render the realized choice a strict improvement under the FSD benchmark applied to perceived lotteries). We then solve the resulting assignment problem with a greedy-refinement heuristic, and run the algorithm with $100$ independent random restarts. We report two rule-level diagnostics. The concentration gain $\phi(f)$ is computed from “full minus one” comparisons (dropping $f$ from $\mathcal{F}$ while keeping the two one-sided-attention defaults. The stability score $\kappa(f)$ is computed by iteratively attempting to delete rules in random orders and recording how often each rule remains necessary to attain the subject’s maximal concentration; we use $500$ random orderings of the library.
@@ Results.
We first report the average level of rule concentration. The average MRCI across subjects is $0.55$ (s.d. $0.05$). This corresponds to an average effective number of rules $N_{eff}\approx 1.85$, suggesting that the average subject behaves as if choosing between fewer than two rules. The density plot associated to $N_{eff}$ is provided in Figure 1.
@@
Finally, three rules are largely inactive by these importance measures, and the interpretation differs across them. For PW (probability weighting) and Id (identity), low $\phi$ and low $\kappa$ do not necessarily mean that these transformations are conceptually irrelevant. They might reflect the fact that these rules, as specified here, primarily re-encode lotteries without reliably generating strict dominance comparisons that would uniquely pin down choices menu-by-menu. In particular, PW distorts probabilities but this distortion alone might not suffice to turn the realized choice into a strict FSD improvement, so it may rarely expand the set of locally admissible assignments. By contrast, DIS (disappointment) appears genuinely non-operative in this dataset under our operationalization: it almost never contributes to maximal concentration and is rarely required to sustain it. We interpret this cautiously as evidence that, within our library and dominance-based discipline, disappointment-based comparisons are not a leading channel for explaining the CPC18 choice patterns.
@@
Finally, we investigate heterogeneity across gender. A simple OLS regression of MRCI on gender (controlling for age) reveals a statistically significant difference: women exhibit an MRCI that is approximately 0.01 points lower than men (see Table A.1 in the Appendix). In terms of the effective number of rules, this implies that men are slightly more concentrated in their rule usage, while women use a slightly more diverse set of rules (higher $N_{eff}$). We decompose this difference in Figure 4. The concentration gains for the dominant rules (MAP and SAL) are slightly higher for men, driving their higher overall MRCI. Conversely, women tend to have higher stability scores across a broader range of secondary rules. This supports the interpretation that male subjects in this sample seem to lock onto a slightly narrower set of rules, whereas female subjects display behavior that is better described by a broader weighted portfolio of rules.
@@ 8 Conclusion
This paper develops a revealed-rule approach to risky choice designed for environments in which perception is plausibly menu-dependent. Rather than taking global rationalizability as the primary source of empirical discipline, we assume that at each menu, the rule used must make the realized choice a strict improvement according to a dominance benchmark applied to the perceived lotteries generated by that rule. Under first-order stochastic dominance, this local discriminability requirement has a strong implication: any rule profile that satisfies it induces a perceived dataset that is automatically cycle-free in revealed-preference terms. Empirical content therefore comes from parsimony across menus; how concentrated a rule-based explanation can be.
@@
The framework is intentionally modular, and has several limitations. First, our clean implication from local admissibility to strong cyclical consistency relies on first-order stochastic dominance; extending the approach to alternative pre-orders may require integrating local discriminability with additional global constraints. Second, we take the rule library as a primitive. A natural next step is to treat the library itself as an empirical object: to compare competing libraries, to penalize library expansion in a principled way, and to study how identification changes as the menu of candidate mechanisms grows. Third, our analysis is nonstructural about the determinants of rule use. Embedding concentration and rule-importance measures in models with cognitive costs, attention constraints, learning from feedback, incentives, or endogenous consideration would clarify when observed parsimony reflects genuine simplicity versus rational adaptation to the choice architecture. More broadly, the approach suggests an agenda in which decision rules are measured and compared across populations and decision environments, providing a bridge between behavioral theory and revealed preference.
Online Appendix
Appendix A Local Discriminability and Cyclical ConsistencyA.1
Definition 8.
[…]
    $x=x_{t}\quad\text{and}\quad y\in A_{t}^{\downarrow},$
    where
[…]
    $x=x_{t}\quad\text{and}\quad y\in A_{t}^{\downarrow\downarrow},$
    where
[…]
  • $R^{\ast}$ denotes the transitive closure of $R$.
[…]
$x\mathrel{R^{\ast}}y\quad\text{and}\quad y\mathrel{P}x.$
[…]
$x\mathrel{R^{\ast}}y\quad\text{and}\quad y\mathrel{R^{\ast}}x.$
Rationality benchmark.
[…]
Proposition A.1.
Proof.
$\widetilde{L}^{i_{t}}_{t,f_{t}}\;>_{\mathrm{FSD}}\;\widetilde{L}^{-i_{t}}_{t,f_{t}}.$
  • For a menu $\widetilde{A}_{t}$, its FSD–downward closure is
$u\,\widetilde{R}^{\ast}\,v\quad\text{and}\quad v\,\widetilde{R}^{\ast}\,u.$
Step 1: Direct revealed preference is FSD–monotone.
We first prove that whenever $x\,\widetilde{R}\,y$ holds, we have
$x\;\geq_{\mathrm{FSD}}\;y.$
$w\;\geq_{\mathrm{FSD}}\;y.$
Case 2: $w=\widetilde{y}_{t}$. By Local Discriminability, we have
$\widetilde{x}_{t}\;>_{\mathrm{FSD}}\;\widetilde{y}_{t},$
$\widetilde{x}_{t}\;\geq_{\mathrm{FSD}}\;y.$
Step 2: Any revealed–preference path is FSD–monotone.
From Step 1, each direct edge $x_{k}\,\widetilde{R}\,x_{k+1}$ implies
$x_{k}\;\geq_{\mathrm{FSD}}\;x_{k+1}.$
hence
$x\;\geq_{\mathrm{FSD}}\;y.$
Thus we have shown:
Step 3: Absence of strong cycles.
Applying the implication from Step 2 twice, we obtain
$u\;\geq_{\mathrm{FSD}}\;v\quad\text{and}\quad v\;\geq_{\mathrm{FSD}}\;u.$
Appendix B Proofs
B.1 Proof of Proposition 2
@@ B.4 Proof of Proposition 5
Appendix C Random Rule Model (RRM)

### v2

@@ 3.3 Examples
$\operatorname{MRCI}(D_{\mathrm{toy}};\geq_{\mathrm{FSD}},\mathcal{F}\setminus\{\textsc{SAL}\})\leq\Big(\tfrac{4}{7}\Big)^{2}+\Big(\tfrac{3}{7}\Big)^{2}\approx 0.51,$
so
$\phi(\textsc{SAL})\ \geq\ \frac{0.76-0.51}{0.76}\ \approx\ 0.33.$
Moreover, SAL appears in every deletion-stable set across permutations, so $\kappa(\textsc{SAL})=1$: it is order-robustly indispensable for sustaining the maximally concentrated explanation. This illustrates the key distinction behind our diagnostics: a rule can be necessary for some menus (like MAP at $t=2$) yet have low gain if it is easily substituted by defaults, whereas a “unifier” like SAL has high gain precisely because it is hard to replicate with any single alternative mechanism.
3.4 Computational Strategy
Computing $\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})$ is a discrete assignment problem: for each menu $t$ we must pick one rule $f_{t}\in\mathcal{F}_{t}^{\mathrm{strict}}$ to maximize the convex concentration objective $\sum_{f\in\mathcal{F}}s_{f}^{2}$. An exact Mixed-Integer Quadratic Programming (MIQP) formulation is given in Appendix E, but because we are maximizing a convex function over a combinatorial feasible set, the problem is NP-hard and the exact solver becomes impractical at the scale of standard experimental datasets.
Our main estimates therefore rely on a fast heuristic detailed in Appendix C. The heuristic exploits a simple implication of convexity: concentration-maximizing explanations tend to load heavily on rules that are high-coverage (strictly admissible at many menus). It proceeds in three phases. (i) Pre-processing: compute the binary admissibility matrix $a_{t,f}=\mathbf{1}\{f\in\mathcal{F}_{t}^{\mathrm{strict}}\}$ once for the subject. (ii) Greedy assignment: for a given ordering of the rules, assign each rule to all currently unassigned menus where it is admissible (moving down the list), which yields a feasible assignment because $\mathcal{F}_{t}^{\mathrm{strict}}$ is never empty. (iii) Random restarts: repeat the greedy pass over many rule orderings (starting with a coverage-first ordering and then random permutations) and retain the assignment with the highest achieved concentration. This routine scales essentially linearly in $T\times|\mathcal{F}|$ per restart and is therefore suitable for subject-by-subject estimation.
We benchmark the heuristic against the exact MIQP on many subsampled instances (Appendix E). The approximation is near-exact: the mean absolute gap is below $10^{-3}$ and more than $96\%$ of estimates fall within a $1\%$ tolerance of the true optimum. These results justify treating the heuristic MRCI as an accurate proxy for maximal attainable concentration in large samples. The same computational routine is also used repeatedly to construct the “full-minus-one” Concentration Gains $\phi$ and the deletion-based Stability Scores $\kappa$ reported in the empirical section.
@@ 4.1 Random Rule Models and two assumptions
  • A probability vector $w=(w_{f})_{f\in\mathcal{F}}\in\Delta^{|\mathcal{F}|-1}$
  • An i.i.d. latent process $(H_{t})_{t\in\mathcal{T}}$ with $\Pr(H_{t}=f)=w_{f}$. Given a menu $A_{t}=\{L_{t}^{1},L_{t}^{2}\}$ and a realized latent rule $H_{t}=f$, the agent chooses the option prescribed by $f$ whenever $f$ yields a strict dominance recommendation at $A_{t}$; when $f$ does not strictly discriminate at $A_{t}$, choice is generated by a fixed fallback that does not depend on the realized menu index. [^8] Any fixed fallback is admissible for the interpretation below. In particular, one may take the fallback to be a randomization between A1 and A2 with fixed weights.
@@ 5 Empirical Application
We apply our framework to the publicly available CPC18 competition dataset (erev_ert_plonsky_2017; plonsky2017_cpc18). [^9] The data can be downloaded on the following website: https://cpc-18.com/data/. Since our rule library and dominance benchmark are defined over lotteries with known probabilities, we restrict attention to the risk problems and exclude ambiguity tasks. The resulting sample comprises $N=686$ subjects (average age 24.9, s.d. 2.93; 54% female) recruited across multiple locations in Israel. Each subject faces $20$ to $28$ distinct binary lottery choice problems, and each problem is repeated for $25$ trials, yielding $500$ to $700$ decisions per subject. The design includes both no-feedback and feedback regimes, [^10] In the feedback regime, repeated exposure and realized outcomes may induce learning or history-dependence. and the choice problems were designed to elicit canonical behavioral anomalies in risky choice, including Allais paradoxes (allais1953; tversky_kahneman1986), certainty-like effects (kahneman_tversky1979), break-even effects (thaler_johnson1990), splitting effects (birnbaum2008), or systematic departures from expected-value reasoning driven by the weighting of rare events (barron_erev2003). For a detailed description of the experimental design, the exact list of problems, and the mapping from problems to targeted phenomena, see erev_ert_plonsky_2017. Summary statistics are provided in Table 2.
@@
  • Notes: The table reports summary statistics for the $N=686$ subjects in the CPC18 dataset. The lower panel reports the Maximum Rule Concentration Index (MRCI), computed subject-by-subject under the rule library $\mathcal{F}$ and the $\geq_{\mathrm{FSD}}$ benchmark. MRCI is estimated with the heuristic algorithm in Section C using $K=500$ independent random restarts per subject; the Effective Number of Rules is $N_{\mathrm{eff}}=1/\operatorname{MRCI}$. The “Share Non-Random” rows report the fraction of subjects for whom we reject the exchangeable random-choice null using the finite-sample permutation test in Section 4. We approximate the exact permutation distribution by Monte Carlo with $B=500$ random permutations. Each permuted-statistic $\operatorname{MRCI}^{(b)}$ is computed using the same heuristic (with $R=100$ restarts).
@@
For each subject $i$, we compute the Maximum Rule Concentration Index $\operatorname{MRCI}(D_{i};\,\geq_{\mathrm{FSD}},\mathcal{F})$ using the optimization routine described in Section C. In practice, we first pre-compute the menu-by-rule admissibility matrix (which records, for each binary problem faced by the subject, which rules in $\mathcal{F}$ render the realized choice a strict improvement under the FSD benchmark applied to perceived lotteries). We then solve the resulting assignment problem with a greedy-refinement heuristic, and run the algorithm with $100$ independent random restarts. We report two rule-level diagnostics. The concentration gain $\phi(f)$ is computed from “full minus one” comparisons (dropping $f$ from $\mathcal{F}$ while keeping the two one-sided-attention defaults). The stability score $\kappa(f)$ is computed by iteratively attempting to delete rules in random orders and recording how often each rule remains necessary to attain the subject’s maximal concentration; we use $100$ random orderings of the library.
@@ Results.
Before turning to concentration, it is useful to document how often each rule can stand alone as a local justification of observed choices. For subject $i$ and rule $f\in\mathcal{F}$, define the (strict) coverage
$\mathrm{Cov}_{i}(f)\;:=\;\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\mathbbm{1}\big\{f\in\mathcal{F}^{\mathrm{strict}}_{it}\big\},\qquad\mathcal{F}^{\mathrm{strict}}_{it}\;:=\;\big\{f\in\mathcal{F}:\ \pi_{f}(x_{it};L^{-}_{it})\;>_{\mathrm{FSD}}\;\pi_{f}(L^{-}_{it};x_{it})\big\},$
where $\mathcal{F}^{\mathrm{strict}}_{it}$ is the set of rules that, when applied to menu $A_{it}=\{x_{it},L^{-}_{it}\}$, render the realized choice a strict FSD improvement in the perceived menu. Coverage is computed before solving the MRCI assignment problem and therefore measures a rule’s mechanical permissiveness: if a rule has near-zero coverage, it cannot matter empirically regardless of its theoretical motivation.
Figure 1 reports mean coverage across subjects. All ten rules have nontrivial coverage in CPC18 (each exceeding roughly $15\%$ on average), including Id and PW. Hence, the importance patterns reported below should not be read as some rules being dormant because they rarely pass the strict-dominance filter; instead, they reflect how strongly a rule helps concentrate admissible assignments once it competes with close substitutes.
Figure 1: Mean rule coverage across subjects. For each subject and rule $f$, coverage is the share of that subject’s observations for which applying $f$ makes the realized choice a strict FSD improvement in the perceived menu. Bars report cross-subject means; rules are ordered as in the baseline library.
We first report the average level of rule concentration. The average MRCI across subjects is $0.545$ (s.d. $0.05$). This corresponds to an average effective number of rules $N_{eff}\approx 1.83$, suggesting that the average subject behaves as if choosing between fewer than two rules. The density plot associated to $N_{eff}$ is provided in Figure 2.
@@
Finally, three rules are largely inactive by these importance measures, but this does not stem from a lack of local admissibility. Figure 1 shows that Id and PW (and, to a lesser extent, DIS) pass the strict-dominance filter in a nontrivial fraction of menus. Their low $\phi$ and low $\kappa$ instead indicate that, conditional on being admissible, these rules are rarely pivotal for attaining maximal concentration: menus that can be assigned to Id or PW can typically be reassigned to higher-coverage “unifiers” (notably MAP and SAL) with little loss in MRCI. In this sense, Id and PW behave as background rules - often locally defensible, but weakly discriminating once other mechanisms are available. DIS is similar: while locally admissible at some menus, it almost never becomes necessary to sustain a high-concentration assignment, suggesting that disappointment-style comparisons are not a leading organizer of CPC18 behavior within our dominance-based discipline.
@@
We then investigate heterogeneity across gender. A simple OLS regression of MRCI on gender (controlling for age) reveals a statistically significant difference: women exhibit an MRCI that is approximately 0.01 points lower than men (see Table A.1 in the Appendix). In terms of the effective number of rules, this implies that men are slightly more concentrated in their rule usage, while women use a slightly more diverse set of rules (higher $N_{eff}$). We decompose this difference in Figure 5. The concentration gains for the dominant rules (MAP and SAL) are slightly higher for men, driving their higher overall MRCI. Conversely, women tend to have higher stability scores across a broader range of secondary rules. This supports the interpretation that male subjects in this sample seem to lock onto a slightly narrower set of rules, whereas female subjects display behavior that is better described by a broader weighted portfolio of rules.
@@ 6 Conclusion
This paper develops a revealed-rule approach to risky choice designed for environments in which perception is plausibly menu-dependent. Rather than taking global rationalizability as the primary source of empirical discipline, we assume that at each menu, the rule used must make the realized choice a strict improvement according to a dominance benchmark applied to the perceived lotteries generated by that rule. Empirical content therefore comes from parsimony across menus; how concentrated a rule-based explanation can be.
@@
The framework is intentionally modular, and has several limitations. First, our implication from local admissibility to strong cyclical consistency relies on first-order stochastic dominance; extending the approach to alternative pre-orders may require integrating local discriminability with additional global constraints. Second, we take the rule library as a primitive. A natural next step is to treat the library itself as an empirical object: to compare competing libraries, to penalize library expansion in a principled way, and to study how identification changes as the menu of candidate mechanisms grows. Third, our analysis is nonstructural about the determinants of rule use. Embedding concentration and rule-importance measures in models with cognitive costs, attention constraints, learning from feedback, incentives, or endogenous consideration would clarify when observed parsimony reflects genuine simplicity versus rational adaptation to the choice architecture. More broadly, the approach suggests an agenda in which decision rules are measured and compared across populations and decision environments, providing a bridge between behavioral theory and revealed preference.
Appendix A Proofs
A.1 Proof of Proposition 2
@@ A.4 Proof of Proposition 5
Appendix B Local Discriminability and Cyclical Consistency
Definition 8.
[…]
    $x=x_{t}\quad\text{and}\quad y\in A_{t}^{\downarrow},$
    where
[…]
    $x=x_{t}\quad\text{and}\quad y\in A_{t}^{\downarrow\downarrow},$
    where
[…]
  • $R^{\ast}$ denotes the transitive closure of $R$.
[…]
$x\mathrel{R^{\ast}}y\quad\text{and}\quad y\mathrel{P}x.$
[…]
$x\mathrel{R^{\ast}}y\quad\text{and}\quad y\mathrel{R^{\ast}}x.$
Rationality benchmark.
[…]
Proposition A.1.
Proof.
$\widetilde{L}^{i_{t}}_{t,f_{t}}\;>_{\mathrm{FSD}}\;\widetilde{L}^{-i_{t}}_{t,f_{t}}.$
  • For a menu $\widetilde{A}_{t}$, its FSD–downward closure is
$u\,\widetilde{R}^{\ast}\,v\quad\text{and}\quad v\,\widetilde{R}^{\ast}\,u.$
Step 1: Direct revealed preference is FSD–monotone.
We first prove that whenever $x\,\widetilde{R}\,y$ holds, we have
$x\;\geq_{\mathrm{FSD}}\;y.$
$w\;\geq_{\mathrm{FSD}}\;y.$
Case 2: $w=\widetilde{y}_{t}$. By Local Discriminability, we have
$\widetilde{x}_{t}\;>_{\mathrm{FSD}}\;\widetilde{y}_{t},$
$\widetilde{x}_{t}\;\geq_{\mathrm{FSD}}\;y.$
Step 2: Any revealed–preference path is FSD–monotone.
From Step 1, each direct edge $x_{k}\,\widetilde{R}\,x_{k+1}$ implies
$x_{k}\;\geq_{\mathrm{FSD}}\;x_{k+1}.$
hence
$x\;\geq_{\mathrm{FSD}}\;y.$
Thus we have shown:
Step 3: Absence of strong cycles.
Applying the implication from Step 2 twice, we obtain
$u\;\geq_{\mathrm{FSD}}\;v\quad\text{and}\quad v\;\geq_{\mathrm{FSD}}\;u.$
Appendix C Computational Strategy
An exact formulation of this problem as a mixed–integer quadratic program (MIQP) is provided in Appendix E. However, maximizing a convex objective function (concentration) over discrete assignments is an NP-hard problem. As demonstrated in our benchmarking exercise (see Appendix E), the MIQP approach scales fast with sample size, rendering it computationally intractable for standard datasets. To address this, we employ a heuristic algorithm that exploits the structure of the objective. Since we are maximizing concentration, optimal solutions by convexity tend to load on rules that can rationalize many menus simultaneously. The algorithm below is designed to efficiently explore such high–coverage, high–concentration assignments.
C.1 Overview of the heuristic
The procedure has three components:
  1. i
  2. $i$
  3. iii
We now describe each component in turn.
Phase 1: Pre–processing
Local admissibility.
$a_{t,f}:=\mathbf{1}\{f\in\mathcal{F}_{t}^{\mathrm{strict}}\},$
Phase 2: Greedy assignment for a fixed rule order
Rule ordering.
$\text{coverage}(f):=\sum_{t=1}^{T}a_{t,f},$
State variables.
Greedy pass.
  1. i
  2. $i$
Phase 3: Random restarts and selection
C.2 Accuracy and Scalability
The heuristic replaces combinatorial optimization with a sequence of fast linear passes. To validate this approach, we benchmarked the heuristic against the MIQP solver on a set of 3,430 subsampled optimization tasks (see Appendix E for full details).
Appendix D Random Rule Model (RRM)